OpenStudy (anonymous):

14(t-2)+7t = 7(3t+4)-10

OpenStudy (anonymous):

14(t-2)+7t = 7(3t+4)-10 (14t - 28) + 7t = (21t + 28) - 10 21t -28 = 21t + 18 -28 = 18 This is absurd, so there might be a problem/mistake in this question

OpenStudy (anonymous):

maybe it 14(t-2)+7 = 7(3t+4)-10